The speeds of an e-scooter (XX km/h) and a pedestrian (YY km/h) measured on a particular stretch of footpath are normally distributed with mean and variance as follows:

  1. ((i))

    State the value of μ\mu and find the value of σ\sigma.

  2. ((ii))

    Given that the speeds of half of the e-scooters measured are found to be within aa km/h of the mean, find aa.

    Find the least value of kk such that Alice is expected to earn a profit.

  3. ((iii))

    A LTA officer stationed himself at the footpath and measured the speeds of 50 e-scooters at random. Find the probability that the 50th e-scooter is the 35th to exceed LTA's legal speed limit of 10 km/h.

  4. ((iv))

    On another day, the LTA officer randomly measured the speeds of 6 e-scooters and 15 pedestrians. Find the probability that the mean speed of the e-scooters is more than twice the mean speed of the pedestrians captured.

  5. ((v))

    Find the probability that the mean speed of nn randomly chosen e-scooters is more than 10 km/h, if nn is large.

Work through it with AI →

It won't just hand you the answer — it asks where you're stuck.How the checker works

Worked solution

10((i))

1. State the value of $\mu$ using the symmetry of the normal distribution.

Since Pr⁡(Y<5.2)=Pr⁡(Y≥7.0)\Pr(Y < 5.2) = \Pr(Y \ge 7.0), by the symmetry of the normal distribution, the mean μ\mu is exactly in the middle of 5.25.2 and 7.07.0.
μ=5.2+7.02=6.1\mu = \frac{5.2 + 7.0}{2} = 6.1

2. Set up an equation involving the standard normal variable $Z$ to find $\sigma$.

Given Pr⁡(Y<5.2)=0.379\Pr(Y < 5.2) = 0.379, we standardize the variable:
Pr⁡(Z<5.2−6.1σ)=0.379\Pr\left(Z < \frac{5.2 - 6.1}{\sigma}\right) = 0.379
−0.9σ≈−0.30808\frac{-0.9}{\sigma} \approx -0.30808

3. Solve for $\sigma$.

σ=−0.9−0.30808≈2.9213\sigma = \frac{-0.9}{-0.30808} \approx 2.9213
σ≈2.92 (to 3 s.f.)\sigma \approx 2.92 \text{ (to 3 s.f.)}

Answer

μ=6.1\mu = 6.1, σ=2.92\sigma = 2.92

10((ii))

1. Translate the 'within a of the mean' condition into a probability equation

X∼N(12.3, 9.9)X \sim \mathrm{N}(12.3,\,9.9) and we require
Pr⁡(12.3−a<X<12.3+a)=0.5.\Pr(12.3 - a < X < 12.3 + a) = 0.5.
By the symmetry of the normal curve, the two tails together carry the remaining 0.50.5, so each tail is 0.250.25:
Pr⁡(X<12.3+a)=0.75.\Pr(X < 12.3 + a) = 0.75.

2. Standardise and solve for a

Using the inverse normal (GC) with mean 12.312.3 and standard deviation 9.9\sqrt{9.9}:
12.3+a=invNorm(0.75, 12.3, 9.9)=14.422…12.3 + a = \mathrm{invNorm}(0.75,\,12.3,\,\sqrt{9.9}) = 14.422\ldots
a=14.422…−12.3=2.122…a = 14.422\ldots - 12.3 = 2.122\ldots
Equivalently, a9.9=0.6745\dfrac{a}{\sqrt{9.9}} = 0.6745, so a=0.6745×3.1464=2.12a = 0.6745 \times 3.1464 = 2.12.
a≈2.12 km/h.a \approx 2.12\text{ km/h}.

Answer

a≈2.12a \approx 2.12 km/h

10((iii))

1. Find the probability that a single randomly chosen e-scooter exceeds 10 km/h.

Let XX be the speed of an e-scooter in km/h. We are given X∼N(12.3,9.9)X \sim \mathrm{N}(12.3, 9.9).
The probability that an e-scooter exceeds 10 km/h is:
p=Pr⁡(X>10)≈0.767605p = \Pr(X > 10) \approx 0.767605

2. Identify the conditions for the 50th e-scooter to be exactly the 35th one to exceed the limit.

For the 50th e-scooter to be the 35th one to exceed the speed limit, two independent events must occur sequentially:

  1. Exactly 34 out of the first 49 e-scooters exceed the speed limit.
  2. The 50th e-scooter exceeds the speed limit.

Let WW be the number of e-scooters out of the first 49 that exceed 10 km/h.
W∼B(49,0.767605)W \sim \mathrm{B}(49, 0.767605)

3. Calculate the final probability.

Required probability=Pr⁡(W=34)×Pr⁡(X>10)\text{Required probability} = \Pr(W = 34) \times \Pr(X > 10)
=[(4934)(0.767605)34(1−0.767605)15]×0.767605= \left[ \binom{49}{34} (0.767605)^{34} (1 - 0.767605)^{15} \right] \times 0.767605
=0.017114×0.767605= 0.017114 \times 0.767605
≈0.013137\approx 0.013137
≈0.0131 (to 3 s.f.)\approx 0.0131 \text{ (to 3 s.f.)}

Answer

0.01310.0131

10((iv))

1. Determine the mean ($\mu$) and variance ($\sigma^2$) of the pedestrian speeds from the information given in part (i).

Pr⁡(Y<5.2)=Pr⁡(Y≥7.0)=0.379\Pr(Y < 5.2) = \Pr(Y \ge 7.0) = 0.379
By the symmetry of the normal distribution:
μ=5.2+7.02=6.1\mu = \frac{5.2 + 7.0}{2} = 6.1
Using the given probability to find σ\sigma:
Pr⁡(Y<5.2)=0.379  ⟹  Pr⁡(Z<5.2−6.1σ)=0.379\Pr(Y < 5.2) = 0.379 \implies \Pr\left(Z < \frac{5.2 - 6.1}{\sigma}\right) = 0.379
Using the inverse normal function on a graphing calculator:
−0.9σ≈−0.308077  ⟹  σ≈2.92134\frac{-0.9}{\sigma} \approx -0.308077 \implies \sigma \approx 2.92134
σ2≈8.53425\sigma^2 \approx 8.53425

2. State the distributions for the sample means $\bar{X}$ and $\bar{Y}$.

For n=6n=6 e-scooters:
Xˉ∼N(12.3,9.96)  ⟹  Xˉ∼N(12.3,1.65)\bar{X} \sim \mathrm{N}\left(12.3, \frac{9.9}{6}\right) \implies \bar{X} \sim \mathrm{N}(12.3, 1.65)
For n=15n=15 pedestrians:
Yˉ∼N(6.1,8.5342515)  ⟹  Yˉ∼N(6.1,0.56895)\bar{Y} \sim \mathrm{N}\left(6.1, \frac{8.53425}{15}\right) \implies \bar{Y} \sim \mathrm{N}(6.1, 0.56895)

3. Formulate the linear combination $W = \bar{X} - 2\bar{Y}$ and calculate its expected value and variance.

Let W=Xˉ−2YˉW = \bar{X} - 2\bar{Y}.
E(W)=E(Xˉ)−2E(Yˉ)=12.3−2(6.1)=0.1\mathrm{E}(W) = \mathrm{E}(\bar{X}) - 2\mathrm{E}(\bar{Y}) = 12.3 - 2(6.1) = 0.1
Var(W)=Var(Xˉ)+22Var(Yˉ)=1.65+4(0.56895)=3.9258\mathrm{Var}(W) = \mathrm{Var}(\bar{X}) + 2^2\mathrm{Var}(\bar{Y}) = 1.65 + 4(0.56895) = 3.9258
Thus, W∼N(0.1,3.9258)W \sim \mathrm{N}(0.1, 3.9258).

4. Calculate the probability that the mean speed of the e-scooters is more than twice the mean speed of the pedestrians.

We need to find Pr⁡(Xˉ>2Yˉ)\Pr(\bar{X} > 2\bar{Y}), which is equivalent to Pr⁡(W>0)\Pr(W > 0).
Using a graphing calculator:
Pr⁡(W>0)≈0.52012\Pr(W > 0) \approx 0.52012
Pr⁡(W>0)=0.520 (to 3 s.f.)\Pr(W > 0) = 0.520 \text{ (to 3 s.f.)}

Answer

0.5200.520

10((v))

1. State the distribution of the sample mean $\bar{X}$.

Since the population is normally distributed, X∼N(12.3,9.9)X \sim \mathrm{N}(12.3, 9.9).
For a sample of size nn, the sample mean Xˉ\bar{X} is exactly normally distributed:
Xˉ∼N(12.3,9.9n)\bar{X} \sim \mathrm{N}\left(12.3, \frac{9.9}{n}\right)

2. Standardise the variable and evaluate the probability as $n$ becomes large.

P(Xˉ>10)=P(Z>10−12.39.9/n)=P(Z>−2.39.9n)\begin{aligned} \mathrm{P}(\bar{X} > 10) &= \mathrm{P}\left(Z > \frac{10 - 12.3}{\sqrt{9.9/n}}\right) \\ &= \mathrm{P}\left(Z > -\frac{2.3}{\sqrt{9.9}} \sqrt{n}\right) \end{aligned}
As n→∞n \to \infty, the term −2.39.9n→−∞-\frac{2.3}{\sqrt{9.9}} \sqrt{n} \to -\infty.
Therefore, P(Z>−∞)=1\mathrm{P}(Z > -\infty) = 1.

Answer

1

AI analysis — may contain mistakes.

Worked this one out on paper? Snap your steps and Gigle's AI math checker will circle the step where it went wrong.

More questions on this topic

  1. 1The owner of Eng Choo Farm claims that the eggs produced at his farm has an average cholesterol content of μ0\mu_0 mg pe…
  2. 2In this question you should state the parameters of any distributions that you use. Crispy Cream Donut Shop sells 2 type…
  3. 3XX is a normal random variable with mean 1 and variance 1, and YY is also a normal random variable with mean μ\mu and…
  4. 4Fishing Company A has two types of fishing vessels, the "Standard" vessel and the "Large" vessel. The amount of fish cau…
  5. 5In this question, you should state clearly all the distributions that you use, together with the values of the appropria…